OpenStudy (anonymous):

what is 3x over 8 divided by 7y over 4

OpenStudy (anonymous):

\[ \frac{\frac{3x}{8}}{\frac{7y}{4}} \] Is that what you are say?

OpenStudy (anonymous):

yes im trying to say

OpenStudy (anonymous):

3x / 14y

OpenStudy (anonymous):